Write the slope-intercept equation for the line that passes through (4, 3) and has slope 5/4. Please show all of your work.
Answer by John10(297)

Use the slope-point formula
y - 3 = 5/4(x - 4)
y - 3 = 5x/4 -5
y = 5x/4 - 2
John10